OpenStudy (anonymous):

A Function Example: Find the linearization, say L(x)=ax+b, of fx=sin2x at the point x=4. Find L(x)

OpenStudy (anonymous):

Could you please include a list of steps with it as well

OpenStudy (anonymous):

Remember y = mx + b? That's all this is. m is your slope. You get that from the derivative of your function f(x) at x = 4. Then use point-slope form to finish. Since you have x = 4, plug that in to f(x) to get y = sin 8 -- that's your "point" of point-slope. You have the slope from f'(x). Plug those into the point-slope form, which is y - y1 = m(x - x1). Done.

OpenStudy (anonymous):

oops sorry i meant to write \[f(x) = \sin ^{2}x\] and x = pi/4